solve using multiplication principle -1/2x=-7/8
Answer by checkley77(12844)

-1/2x=-7/8
-x/2=-7/8
-8x=-14
x=-14/-8
x=7/4 ans.
Proof:
-1/2*7/4=-7/8
-7/8=-7/8